OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 –3x + 2y = 1 Answer A)there is no y value as there is no solution to the system. B)4 C)2 D)y can be any number as there are infinitely many solutions to the system.

OpenStudy (anonymous):

3x + y = 11 -3x + 2y = 1 add the two: 3x - 3x + y + 2y = 11 + 1 3y = 12 y = 4 B.